OpenStudy (anonymous):

i need alot of help i have a paper of promblems n i have no clue how to do them do any of you know how to btw im in 9th grade

OpenStudy (anonymous):

whats the first problem?

OpenStudy (anonymous):

5x + 3 =23

OpenStudy (anonymous):

a muti-step equation

OpenStudy (anonymous):

ya i know

OpenStudy (anonymous):

Okay, the first thing you need to do is subtract 3 from each side. Your new equation would be 5x= 23-3 so 5x = 20. Now just divide each side by 5. and there is your answer

OpenStudy (anonymous):

thans bubbles can u help me on some more

OpenStudy (anonymous):

yep yep! :)

OpenStudy (anonymous):

k here comes another 1

OpenStudy (anonymous):

4 = 3a -14

OpenStudy (anonymous):

Okay. It is exactly like the other problem with different numbers. I want you to learn, so I am going to ask you a question. What did I do first on the first equation?

OpenStudy (anonymous):

u subtracted 3 from both sides

OpenStudy (anonymous):

Great. This time we want to combine the like numbers again (which is what we did when we subtracted the 3 from the 23) so we will add 14 to both sides for this equation. This will give you 14 + 4 = 3a. Do you know what to do next?

OpenStudy (anonymous):

add 14 n 4 n get 18 n divide by 3

OpenStudy (anonymous):

Perfect! :) any more problems?